How would I solve this problem:
b+b+18=4b

Answer by Maths68(1474)

b+b+18=4b
2b+18=4b
2b=4b-18
2b-4b=-18
-2b=-18
Divide by -2 both sides of above equation
-2b/-2=-18/-2
b=9